1 For the hypothesis test H0 mu<=5 against H1: mu >5 with variance unknown and n=11, find the best approximation for the P-value for the test statistic t0=1.945.

0.25 ≤ p ≤ 0.50

0.10 ≤ p ≤ 0.25

0.010 ≤ p ≤ 0.025

0.025 ≤ p ≤ 0.050

0.0025 ≤ p ≤ 0.0050

2

The probability of type II error increases if the difference between the hypothesized values of the parameter increases, assuming that the sample size and other parameters do not change.

True

False

3

A random sample of 450 registered voters in Phoenix is asked if they favor the use of oxygenated fuels year-round to reduce air pollution. If more than 360 voters respond positively,
Can we conclude that more than 60% of the voters favor the use of these fuels. Use Alpha=0.05.
       Yes
       Not able to decide
       No
       Missing information
Answer #1

i p value = 1-p(tio {to) = 0402 (1025 <p 2.050 1 Fourth option tue. 2) If the difference increases, power increases 4 hencefor query in above, comment.